A scuba diver measures an increase in pressure of around 105 Pa
upon descending by 10 m, what is the change in force per square
centimetre on the diver’s body?
1
Expert's answer
2020-05-04T12:47:37-0400
"\frac{\Delta F}{A}=10^5(10^{-2})^2=10\frac{N}{cm^2}"